# The n_TOF Collaboration

### v1

@@ III.6 Systematic uncertainty
IV Resonance analysis
The resonances in the capture yield were analyzed with the multilevel multichannel R-matrix code SAMMY [42]. The code allows for a Bayesian fitting of the capture yield in the Resolved Resonance Region (RRR) using the Reich-Moore R-matrix approximation [52], including several experimental effects like the Doppler and RF broadening, the multi-isotopic composition of the samples and the self-shielding and multiple-scattering effects. With SAMMY it is also possible to include the residual background present in the yield (see Sec. III.2) by fitting the valleys between resonances.
IV.1 Resonance analysis with SAMMY
One of the main goals of this measurement was to minimize the multiple-scattering effects in the capture yield. For this purpose, two samples with different thickness were used for each isotope. The region with the wide s-wave resonances between 1 and 10 keV was analyzed using the very thin samples, where multiple-scattering effects are much less relevant. The rest up to 100 keV was analyzed with the thick samples.
@@ IV.1 Resonance analysis with SAMMY
A total of 33 resonances have been observed and analyzed for 50Cr and 51 for 53Cr. The complete list of resonance parameters (with their correlation when applicable) and radiative kernels (see Sec. IV.3) are listed in Tables 6 and 7 of the Appendix.
IV.2 Results and discussion between 1 and 10 keV
The range between 1 and 10 keV is the most important for criticality benchmarks because of a cluster of strong s-wave resonances in both 50Cr and 53Cr. These resonances are also the main source of multiple-scattering effects, which severely affected previous measurements. The measured capture yield with the corresponding SAMMY fit are displayed and compared with the expected yields from the JEFF-3.3, CENDL-3.2 and INDEN [10] evaluations in the top panels of Fig. 8. The figure illustrates very clearly the large difference between evaluations. In this particular region, our result for 50Cr is in very good agreement with JEFF-3.3 but significantly deviates from the other evaluations. Note that although the shape of the resonance is very different from CENDL-3.2, the resonance parameters are such that the resonance kernel (see Sec. IV.4) is in very good agreement with this evaluation. Regarding 53Cr, the best agreement, but surely not perfect, is reached with CENDL-3.2. Notably, for both isotopes the INDEN evaluation clearly overestimates the cross sections. The structures at 1.6 keV and 2.6 keV correspond to the first resonances of 52Cr and 63Cu impurities.
@@ IV.2 Results and discussion between 1 and 10 keV
To illustrate the problems related to the use of thick samples in previous measurements (significantly thicker than our thick targets), the resonance parameters from the analysis of the thin samples has been used in SAMMY to predict the yield corresponding to our thick samples. As shown at the middle panels of Fig. 8, the predictions do not reproduce our spectra with the parameters deduced from our thin sample, specially for the 53Cr case. The total capture yield can be expressed at the sum of individual components $Y_{i}$ as $Y(E_{n})=\sum_{i}Y_{i}(E_{n})$, where the index $i$ indicates the number of scatterings before the capture. To better visualize the importance of the multiple-scattering effects, we display in the bottom panels of Fig. 8 the fraction of the total capture yield that corresponds to $Y_{0}$, i.e., the fraction of captures without any previous scattering. It is clear that the multiple-scattering components are an important contribution to the capture yield in our thick samples, contributing in some regions by more than 50% to the yield. For illustration, the figure also shows $Y_{0}$ (based on our parameters from the thin samples) that would correspond to the samples used in the previous experiments [5, 6, 7, 8, 9]. These samples were 8 to 30 and 7 to 12 times thicker than our thin 50Cr and 53Cr samples, respectively.
IV.3 Results and discussion between 10 and 100 keV
@@ IV.4 Radiative kernels
IV.5 Integral cross sections (MACS)
@@
where $E_{n}$ is the neutron energy in the center-of-mass system and $\sigma(E_{n})$ is the point-wise capture cross section. Since the measurement described herein only covers the region below $E_{n}=100~{}$keV, we have used JEFF-3.3 and CENDL-3.2 to extrapolate the measured cross section up to 300 keV and evaluate the difference in the MACS when using one or another evaluation for the extrapolation.
@@
  • Regarding 50Cr, our $MACS$ is in clear disagreement (30% lower) with the recent INDEN evaluation, and 8-9% lower, slightly beyond one standard deviation, than JEFF-3.3 and CENDL-3.2. Our value is then in a remarkable agreement within uncertainties with the $MACS$ value of Pérez-Maroto et al. obtained recently by activation at the CNA HiSPANoS facility.
@@
V Summary and conclusions
A neutron capture time-of-flight measurement of 50Cr and 53Cr has been successfully performed at the n_TOF-EAR1 facility of CERN. The capture set-up consisted on four C6D6 Total Energy detectors characterized by their very low neutron sensitivity. Two highly enriched chromium oxide samples were used for each isotope: a very thin one was used to analyze the yield between 1 and 10 keV, and a thicker one for the rest of the neutron energy range up to 100 keV. All samples were thinner than any of those used in previous measurements, thus strongly suppressing the multiple-scattering effects that affected the previous measurements, thus causing the discrepancies present in the evaluated cross section libraries.
The capture yield has been measured from 1 to 100 keV with an accuracy between 5 and 9%, fulfilling the requirements of the NEA HPRL request. A total of 33 resonances of 50Cr and 51 resonances of 53Cr have been identified and analysed with SAMMY, resulting in a new set of resonance parameters. Overall, the comparison of radiative kernels indicates a 50Cr capture cross section $\sim$7lower than JEFF-3.3 (and thus ENDF/B-VIII.0, JENDL-5) and CENDL-3.2, and almost 40% lower than INDEN (and thus ENDF/B-VIII.1). Very similar difference is also reached when the comparison is made for the $\text{MACS}_{30}$ integral cross section. On the other hand, the MACS extracted from the 50Cr n_TOF differential cross section measurement is in an excellent agreement with that obtained recently by neutron activation at CNA HiSPANoS [21].
The discrepancies with the evaluations are much larger in the case of 53Cr. According to the radiative kernels, our cross section is on average 35% larger than JEFF-3.3 (and thus ENDF/B-VIII.0, JENDL-5), 10% lower than CENDL-3.2 and a remarkable 60% lower than INDEN (and thus ENDF/B-VIII.1). When comparing the $\text{MACS}_{30}$ values, a very good agreement if found with CENDL-3.2, while our result is 70% larger than the one expected from INDEN. Accordingly, the sizeable increase in the chromium cross sections proposed in the recent INDEN evaluation is not supported by the new experimental data, which calls for a re-evaluation of cross section including the data presented herein.
Acknowledgements
This measurement has received funding from the Euroatom research and training programme 2014-2018 under grant agreement No 847594 (ARIEL), and from the Spanish national projects RTI2018-098117-B-C21, PID2019-104714GB-C22, PID2021-123879OB-C21 and PID2022-138297NB-C21.
@@ Acknowledgements
Appendix (Resonance parameters and kernels)
Table 6: Resonance parameters of 50Cr obtained in this work. The uncertainty of the resonance energy has been considered as 0.1% based on time-to-energy calibration. The uncertainty of the resonance widths is the one obtained in the SAMMY fitting. If a value has no uncertainty it means that it has not been fitted but adopted from an evaluated library (see Section IV.3).
@@
Table 7: Resonance parameters of 53Cr obtained in this work. The uncertainty of the resonance energy has been considered as 0.1% based on time-to-energy calibration. The uncertainty of the resonance widths is the one obtained in the SAMMY fitting. If a value has no uncertainty it means that it has not been fitted but adopted from an evaluated library (see Section IV.3).

### v2

@@ 3.6 Systematic uncertainty
[…]
4 Resonance analysis
The resonances in the capture yield were analysed with the multilevel multichannel R-matrix code SAMMY [42]. The code allows for a Bayesian fitting of the capture yield in the Resolved Resonance Region (RRR) using the Reich-Moore R-matrix approximation [53], including several experimental effects like the Doppler and RF broadening, the multi-isotopic composition of the samples and the self-shielding and multiple-scattering effects. With SAMMY it is also possible to include the residual background present in the yield (see Sec. 3.2) by fitting the valleys between resonances.
4.1 Resonance analysis with SAMMY
One of the main goals of this measurement was to minimize the multiple-scattering effects in the capture yield. For this purpose, two samples with different thickness were used for each isotope. The region with the wide s-wave resonances between 1 and 10 keV was analysed using the very thin samples, so multiple-scattering effects are much less relevant. The rest up to 100 keV was analysed with the thick samples.
@@ 4.1 Resonance analysis with SAMMY
A total of 33 resonances have been observed and analysed for 50Cr and 51 for 53Cr. The complete list of resonance parameters (with their correlation when applicable) and radiative kernels (see Sec. 4.3) are listed in Tables A and A of the Appendix.
4.2 Results and discussion between 1 and 10 keV
The range between 1 and 10 keV is the most important for criticality benchmarks because of a cluster of strong s-wave resonances in both 50Cr and 53Cr. These resonances are also the main source of multiple-scattering effects, which severely affected previous measurements. The measured capture yield with the corresponding SAMMY fit are displayed and compared with the expected yields from the JEFF-3.3, CENDL-3.2 and INDEN [10] evaluations in the top panels of Fig. 8. The figure illustrates very clearly the large difference between evaluations. In this particular region, our result for 50Cr is in very good agreement with JEFF-3.3 but significantly deviates from the other evaluations. Note that although the shape of the resonance is very different from CENDL-3.2, the resonance parameters are such that the resonance kernel (see Sec. 4.4) is in very good agreement with this evaluation. Regarding 53Cr, the best agreement, but surely not perfect, is reached with CENDL-3.2. Notably, for both isotopes the INDEN evaluation clearly overestimates the cross sections. The structures at 1.6 keV and 2.6 keV correspond to the first resonances of 52Cr and 63Cu impurities.
@@ 4.2 Results and discussion between 1 and 10 keV
To illustrate the problems related to the use of thick samples in previous measurements (significantly thicker than all of our samples), the resonance parameters from the analysis of the thin samples has been used in SAMMY to predict the yield corresponding to our thick samples. As shown at the middle panels of Fig. 8, the predictions do not reproduce our spectra with the parameters deduced from our thin sample, specially for the 53Cr case, evidencing the limited capability of SAMMY for estimating accurately the multiple-scattering interactions prior to the neutron capture for very thick samples. The total capture yield can be expressed at the sum of individual components $Y_{i}$ as $Y(E_{n})=\sum_{i}Y_{i}(E_{n})$, where the index $i$ indicates the number of scatterings before the capture. To better visualize the importance of the multiple-scattering effects, we display in the bottom panels of Fig. 8 the fraction of the total capture yield that corresponds to $Y_{0}$, i.e., the fraction of captures without any previous scattering. It is clear that the multiple-scattering components are an important contribution to the capture yield in our thick samples, contributing in some regions by more than 50% to the yield. For illustration, the figure also shows $Y_{0}$ (based on our parameters from the thin samples) that would correspond to the samples used in the previous experiments [5, 6, 7, 8, 9]. These samples were 8 to 30 and 7 to 12 times thicker than our thin 50Cr and 53Cr samples, respectively.
4.3 Results and discussion between 10 and 100 keV
@@ 4.4 Radiative kernels
4.5 Integral cross sections (MACS)
@@
where $E_{n}$ is the neutron energy in the centre-of-mass system and $\sigma(E_{n})$ is the point-wise capture cross section. Since the measurement described herein only covers the region below $E_{n}=100~$keV, we have used JEFF-3.3 and CENDL-3.2 to extrapolate the measured cross section up to 300 keV and evaluate the difference in the MACS when using one or another evaluation for the extrapolation.
@@
  • Regarding 50Cr, our $\text{MACS}_{30}$ is in clear disagreement (30% lower) with the recent INDEN evaluation, and 8-9% lower, slightly beyond one standard deviation, than JEFF-3.3 and CENDL-3.2. Our value is then in a remarkable agreement within uncertainties with the $\text{MACS}_{30}$ value of Pérez-Maroto et al. obtained recently by activation at the CNA HiSPANoS facility.
@@
[…]
5 Summary and conclusions
A neutron capture time-of-flight measurement of 50Cr and 53Cr has been successfully performed at the n_TOF-EAR1 facility of CERN. The capture set-up consisted on four C6D6 Total Energy detectors characterized by their very low neutron sensitivity. Two highly enriched chromium oxide samples were used for each isotope: a very thin one was used to analyse the yield between 1 and 10 keV, and a thicker one for the rest of the neutron energy range up to 100 keV. All samples were thinner than any of those used in previous measurements, thus strongly suppressing the multiple-scattering effects that affected the previous measurements causing the discrepancies present in the evaluated cross section libraries.
The capture yield has been measured from 1 to 100 keV with an accuracy between 5 and 9%, fulfilling the requirements of the NEA HPRL request. A total of 33 resonances of 50Cr and 51 resonances of 53Cr have been identified and analysed with SAMMY, resulting in a new set of resonance parameters. Overall, the comparison of radiative kernels indicates a 50Cr capture cross section $\sim$7lower than JEFF-3.3 (and thus ENDF/B-VIII.0, JENDL-5) and CENDL-3.2, and almost 40% lower than INDEN (and thus ENDF/B-VIII.1). Very similar difference is also reached when the comparison is made for the $\text{MACS}_{30}$ integral cross section. On the other hand, the $\text{MACS}_{30}$ extracted from the 50Cr n_TOF differential cross section measurement is in an excellent agreement with that obtained recently by neutron activation at CNA HiSPANoS [21].
The discrepancies with the evaluations are much larger in the case of 53Cr. According to the radiative kernels, our cross section is on average 35% larger than JEFF-3.3 (and thus ENDF/B-VIII.0, JENDL-5), 10% lower than CENDL-3.2 and a remarkable 60% lower than INDEN (and thus ENDF/B-VIII.1). When comparing the $\text{MACS}_{30}$ values, a very good agreement if found with CENDL-3.2, while our result is 70% larger than the one expected from INDEN.
The sizeable increase in the chromium cross sections proposed in the recent INDEN evaluation is not supported by the results shown herein. Therefore, a re-evaluation of these cross sections using the new experimental data might contribute to solve the actual issues with the criticality safety benchmarks sensitive to chromium and stainless steel.
[…]
Acknowledgements This measurement has received funding from the Euroatom research and training programme 2014-2018 under grant agreement No 847594 (ARIEL), and from the Spanish national projects RTI2018-098117-B-C21, PID2019-104714GB-C22, PID2021-123879OB-C21 and PID2022-138297NB-C21.
@@ 5 Summary and conclusions
Appendix A Resonance parameters and kernels
Table 6: Resonance parameters of 50Cr obtained in this work. The uncertainty of the resonance energy has been considered as 0.1% based on time-to-energy calibration. The uncertainty of the resonance widths is the one obtained in the SAMMY fitting. If a value has no uncertainty it means that it has not been fitted but adopted from an evaluated library (see Section 4.3).
@@ Appendix A Resonance parameters and kernels
Table 7: Resonance parameters of 53Cr obtained in this work. The uncertainty of the resonance energy has been considered as 0.1% based on time-to-energy calibration. The uncertainty of the resonance widths is the one obtained in the SAMMY fitting. If a value has no uncertainty it means that it has not been fitted but adopted from an evaluated library (see Section 4.3).